# Human Creativity in the Age of LLMs

### v1

@@ 5.3. Limitations & Future Work
Firstly, measuring creativity itself presents challenges. Though we employed both divergent and convergent thinking tasks (AUT and RAT) to capture a broad spectrum of creative processes, these tasks were limited to verbal responses and conducted over short time periods. Creativity, however, is a much more complex and nuanced phenomenon. Future work should expand the battery of tasks to include more natural, real-world creative activities, such as writing advertisements or solving complex problems. Additionally, creativity tasks that are non-verbal or visual, like those supported by diffusion models, could be explored. Experiments using visual creativity tasks, such as the Test of Creative Thinking Drawing Production (TCT-DP) or the Evaluation of Potential Creativity (EoPC) (Lubart, 2016), would provide deeper insights into how LLMs impact creativity in non-text domains.
@@
Our study also focused exclusively on the active, conscious aspects of creativity, largely due to the controlled lab environment. Yet, creativity often involves unconscious processes that are harder to capture in such settings. Another limitation lies in the short exposure period in our lab-based study. In real-world settings, exposure to LLMs is often prolonged, integrated into daily workflows, and thus, likely produces more significant and lasting effects on creativity. In contrast, our study’s shorter exposure periods may have resulted in smaller effect sizes, limiting our ability to fully understand the long-term implications of LLM use. Future research should explore prolonged interactions with LLMs to better reflect real-world scenarios and study potential cumulative impacts on creativity.
@@
Additionally, the design of our LLM-based guidance presents limitations. In the convergent thinking task, the guidance provided was naturally more verbose than direct answers, which may have inadvertently influenced the results. Future studies should control for verbosity and ensure that the nature of guidance is standardized to more accurately assess the impact of LLM “coaching” versus direct assistance. Finally, although this study represents an early attempt to experimentally measure the impact of LLMs on human creativity, the static nature of our LLM interaction may not fully reflect real-world applications. In practical use cases, AI tools often engage in dynamic, interactive exchanges where users can refine inputs, seek clarifications, or adjust outputs. Future research should investigate how conversational and adaptive LLMs influence creativity over time, as these systems more closely resemble how people use AI in real-world creative processes.
@@ 6. Conclusion
Through this work, we sought to understand the impact of LLMs on human creativity. We conducted two parallel experiments on divergent and convergent thinking, two key components of creative thinking. Taken together, these experiments shed light on the complex relationship between human creativity and LLM assistance, suggesting that while AI can augment creativity, the mode of assistance matters greatly and can shape long-term creative abilities. In closing, we hope this work offers a template to experimentally evaluate the impact of generative AI on human cognition and creativity.
@@ Acknowledgements.
We thank Jake Hofman and Polly Denny for helpful discussions on this topic.
@@ A.1.1. LLM Model Specification
A.1.2. Prompt Design
”What are some creative uses for a [OBJECT]? The goal is to come up with creative ideas, which are ideas that strike people as clever, unusual, interesting, uncommon, humorous, innovative, or different. List creative uses for a [OBJECT].”
@@ A.1.3. System Prompt
“List of Strategies” condition
@@ A.2. Sample LLM Responses
Figure 12. Sample LLM responses for the exposure rounds.
@@ B.2. Sample LLM Responses
Figure 14. Sample LLM responses for the exposure rounds.

### v2

@@ 5.3. Limitations & Future Work
Firstly, measuring creativity itself presents challenges. Though we employed both divergent and convergent thinking tasks (AUT and RAT) to capture a broad spectrum of creative processes, these tasks were limited to verbal responses and conducted over short time periods. Creativity, however, is a much more complex and nuanced phenomenon. Future work should expand the battery of tasks to include more natural, real-world creative activities, such as writing advertisements, designing magic tricks, or solving complex problems. Additionally, creativity tasks that are non-verbal or visual, like those supported by diffusion models, could be explored. Experiments using visual creativity tasks, such as the Test of Creative Thinking Drawing Production (TCT-DP) or the Evaluation of Potential Creativity (EoPC) (Lubart, 2016), would provide deeper insights into how LLMs impact creativity in non-text domains.
@@
Our study also focused exclusively on the active, conscious aspects of creativity, largely due to the controlled lab environment. Yet, creativity often involves unconscious processes that are harder to capture in such settings. Another limitation lies in the short exposure period in our lab-based study. In real-world settings, exposure to LLMs is often prolonged, integrated into daily workflows, and thus, likely produces more significant and lasting effects on creativity. In contrast, our study’s shorter exposure periods may have resulted in smaller effect sizes, limiting our ability to fully understand the long-term implications of LLM use. Future research should explore prolonged interactions with LLMs to better reflect real-world scenarios and study potential cumulative impacts on creativity. We acknowledge that limiting the AUT objects in Experiment-1 to five is a limitation. We selected them due to the strong correlation between our automated scoring method and human judgments for these objects, which enabled us to evaluate  10,000 ideas at scale without prohibitive human annotation, in line with other large-scale creativity studies that have operated under similar constraints (Ashkinaze et al., 2024).
@@
Additionally, the design of our LLM-based guidance presents limitations. In the convergent thinking task, the guidance provided was naturally more verbose than direct answers, which may have inadvertently influenced the results. The list-of-strategies and guidance conditions introduced more cognitive load and hence the comparisons with the default conditions during exposure may have been unfair. Future studies should control for verbosity and ensure that the nature of guidance is standardized to more accurately assess the impact of LLM “coaching” versus direct assistance. However, the additional cognitive load during exposure may be beneficial for human cognition long-term as has been shown in education-related experiments. Another approach could be to measure the human performance in the final 30 seconds of the AUT task as the participants may have completed reading/processing LLM output by this time. Future work can also explore the trends in performance split by the different rounds of the experiment. Moreover, participants in the coach conditions may have performed worse during the test due to fatigue build-up in the exposure rounds. Or it could be due to the increased dependency on LLM responses. Future work should investigate the exact mechanisms behind these effects. Finally, although this study represents an early attempt to experimentally measure the impact of LLMs on human creativity, the static nature of our LLM interaction may not fully reflect real-world applications. In practical use cases, AI tools often engage in dynamic, interactive exchanges where users can refine inputs, seek clarifications, or adjust outputs. Future research should investigate how conversational and adaptive LLMs influence creativity over time, as these systems more closely resemble how people use AI in real-world creative processes.
@@ 6. Conclusion
Through this work, we sought to understand the impact of LLMs on human creativity. Our work moves beyond a single-use scenario or model-only evaluations. We conducted two parallel experiments on divergent and convergent thinking, two key components of creative thinking. By exploring repeated exposure, we attempt to approximate real-world interactions with LLMs and investigate the lasting impacts on human creative processes. Taken together, these experiments shed light on the complex relationship between human creativity and LLM assistance, suggesting that while AI can augment creativity, the mode of assistance matters greatly and can shape long-term creative abilities. In closing, we hope this work offers a template to experimentally evaluate the impact of generative AI on human cognition and creativity.
@@ Acknowledgements.
We are grateful to Polly Denny for introducing us to philosophical and historical perspectives on creativity. We are grateful to Jake Hofman for his feedback on our experimental design, and to him, Dan Goldstein, and David Rothschild collectively for their sports metaphor on human–AI relationships, which inspired much of this work. We also appreciate Hope Schroeder for recommending pivotal GenAI+creativity papers, and Jessica Bo, Lilio Mok, and members of the CSSLab at the University of Toronto for their valuable discussions leading up to the experiments. Finally, we thank Peter Organisciak and his group for generously providing access to their automated scoring API for AUT ideas.
@@ A.1.1. Model Specification
A.1.2. Input Prompt: ”What are some creative uses for a [OBJECT]? The goal is to come up with creative ideas, which are ideas that strike people as clever, unusual, interesting, uncommon, humorous, innovative, or different. List creative uses for a [OBJECT].”
@@ A.1.3. System Prompt
“List of Strategies” condition (Figure 11)
@@ A.2. Sample LLM Responses
Table 1 shows the sample LLM responses in Experiment-1.
AUT Object LLM Ideas LLM Strategies
Tire Garden Planter: Cut a tire in half and use it as a unique planter for flowers or vegetables.
Swing: Hang a tire from a tree to create a fun outdoor swing for kids.
Coffee Table: Stack a few tires on top of each other, paint them, and add a glass top to create a unique coffee table.
Sandbox: Fill a large tire with sand to create a small sandbox for children.
Tire Ottoman: Stuff a tire with cushioning material and cover it with fabric to create a unique ottoman.
Tire Pond: Dig a hole in your backyard, place a tire in it, and fill it with water to create a small pond.
Tire Climbing Wall: Attach several tires to a wall to create a climbing wall for kids.
What can you replace on the tire to improve it or give it a new use?
How can you combine the tire with other items or ideas to create something new?
How can the tire be adapted or modified to serve another purpose?
How can altering the shape, size, or color of the tire create new uses or improvements?
What completely different uses can you brainstorm for the tire, beyond its original purpose?
What might happen if parts of the tire were removed or simplified? How does this change its potential uses?
What would happen if you reversed the tire’s purpose or use? Would it function differently or serve a new role?”
Pants Plant Pot Cover: Use pants to cover unsightly plant pots. The pockets can also hold gardening tools.
Art Canvas: Use the fabric of the pants as a canvas for painting or drawing.
Hammock for Small Pets: Sew the ends of the legs together to create a cozy hammock for small pets.
Wine Bottle Bag: Use a pant leg as a chic, recycled wine bottle gift bag.
Patchwork Quilt: Combine various pants to create a unique patchwork quilt.
DIY Bag: Convert an old pair of jeans into a stylish tote bag.
Halloween Decorations: Stuff the pants with hay or newspapers and pair with a shirt to make a scarecrow.
What could you replace the fabric or parts of the pants with to create something new?
How can you combine pants with another item to make a unique product?
How could you modify pants to serve a completely different purpose?
What happens if you change the shape, size, or color of the pants?
What entirely different function could pants serve?
What if you remove certain parts of the pants? How does this change their use?
What if you reverse the way pants are used or worn? What new function could they serve?
Table 1. Sample LLM responses for the exposure rounds in the divergent thinking experiment.
@@ B.2. Sample LLM Responses
Remote Associates Task LLM Answer LLM Guidance
flake, mobile, cone (snow) Snow (Snowflake, Snowmobile, Snowcone) Possible associated words for ”flake” are snow, cereal, skin, coconut, and pastry. Possible associated words for ”mobile” are phone, sculpture, home, device, and car. Possible associated words for ”cone” are ice cream, volcano, traffic, pine, and speaker. Now, try to identify any connections between these words and see if you can come up with a fourth word that is associated with all three. It might be helpful to jot down the words on paper to visually see any connections. Good luck!
dream, light, break (day) Day (Daydream, Daybreak, Daylight) Possible associated words for ”dream” are sleep, subconscious, nightmare, fantasy, and aspiration. Possible associated words for ”break” are fracture, pause, vacation, interruption, and opportunity. Possible associated words for ”light” are lamp, bright, illuminate, candle, and photon. Take a moment to think about any connections between these words and try to come up with a fourth word that is associated with all three. Remember to jot down the words on paper if it helps you visualize any connections. Keep up the good work!
aid, rubber, wagon (band) Band (Bandaid, Rubber Band, Bandwagon) Possible associated words for ”aid” are help, support, assist, rescue, first. Possible associated words for ”wagon” are cart, horse, transport, carry, goods. Possible associated words for ”rubber” are elastic, tire, eraser, latex, ball. Now, can you identify any connections between these words? Is there a fourth word that could connect them? Remember, it might be helpful to jot down the words on paper and see if any connections jump out at you.
Table 2. Sample LLM responses for the exposure rounds in the convergent thinking experiment.